OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)? fan and medal

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

@TheRealMeeeee

OpenStudy (fibonaccichick666):

any ideas on how to begin?

OpenStudy (anonymous):

well i think its (y + 2) = -2(x + 5)

OpenStudy (fibonaccichick666):

how did you get the -2?

OpenStudy (anonymous):

becuase point slope formula is y-y=mx(x-x1) and the points (-2,-5) woul make the signs turn into postives

OpenStudy (fibonaccichick666):

i meant the red two \[(y + 2) = \color{red}{-2}(x + 5)\]

OpenStudy (fibonaccichick666):

to start, but point slope, you have almost correct for the formula, it's not mx, it's just m

OpenStudy (anonymous):

this is my question

OpenStudy (fibonaccichick666):

I just want you to tell me why

OpenStudy (anonymous):

(-2,-5) and (2,3) first we need slope m=y2-y1/x2-x1 m=3+5/2+2 m=8/4 m=2. so far we have y=2x+b now sub in any point. i will use (2,3) where 2=x and 3=y 3=2(2)+b 3=4+b -1=b now we have our equaion. y=2x-1

OpenStudy (anonymous):

those are my points

OpenStudy (fibonaccichick666):

of course... you are guessing. I should have known. You shouldn't do that. It makes those helping you think you know something you do not. So, to start, what does the m represent?

OpenStudy (anonymous):

oh sorry. i didn't do point slope form, idk what that is?

OpenStudy (anonymous):

no im not guessing im trying to plug in the points uhh duuhh

OpenStudy (anonymous):

this website should help. our teachers didn't teach us it, since they felt we had to go at a quicker pace but heres a website to help you http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

okay its not -2 its either b or c

OpenStudy (anonymous):

nvm its c i got the equation